# Learning 3D Representations from Procedural 3D Programs

### v1

@@ 4.1 Object Classification
OcCo [27] 92.1 84.85 85.54 78.79
Point-BERT [42] 93.2 87.43 88.12 83.07
Scratch [42] 91.4 79.86 80.55 77.24
Point-MAE-SN [19] 93.8 90.02 88.29 85.18
Point-MAE-Zero 93.0 90.36 88.64 85.46
@@
Tab. 1 reports the object classification results in the transfer learning setting. On ModelNet40, Point-MAE-Zero achieves performance comparable to Point-MAE-SN. We attribute the performance gap to the substantial domain difference between the clean 3D shapes in ModelNet40 and the procedurally synthesized 3D models. In contrast, the domain gap between ShapeNet and ModelNet40 is smaller, resulting in closer performance between models. On ScanObjectNN, Point-MAE-Zero outperforms Point-MAE-SN across all three variants. This underscores the advantages of pretraining on the procedurally synthesized 3D models which encompass diverse geometry and topology. Both Point-MAE-Zero and Point-MAE-SN exceed the performance of models trained from scratch and previous methods that used alternative self-supervised learning approaches (e.g., OcCo [27] or Point-BERT [42]).
@@
Table 2: Few-shot classification on ModelNet40. We evaluate performance on four $n$-way, $m$-shot configurations. For example, 5w/10s denotes a 5-way, 10-shot classification task. The table reports the mean classification accuracy (%) and standard deviation across 10 independent runs for each configuration. Top: Results from existing methods for comparison. Bottom: Comparison with our baseline methods.
@@
We conduct few-shot classification on ModelNet40 using a standard $n$-way, $m$-shot configuration, where $n$ denotes the number of randomly chosen classes, and $m$ represents the number of instances sampled per class. For evaluation, 20 previously unseen samples are randomly selected from each of the $n$ classes. We perform 10 independent runs and report the average accuracy with standard deviation for each configuration.
Table 2 shows the results for $n=\{5,10\}$ and $m=\{10,20\}$. Our observations are consistent with those in the transfer learning experiments — Point-MAE-Zero performs slightly below Point-MAE-SN on ModelNet40. Both Point-MAE-Zero and Point-MAE-SN notably outperform models trained from scratch, as well as previous methods that utilize different network architectures (e.g., DGCNN [28]) or pretraining strategies.
@@ 4.2 Part Segmentation
The 3D part segmentation task aims to predict a part class label for each point within an object. We evaluate our model and baseline methods on the ShapeNetPart dataset [41], which includes 16,881 models across 16 categories. In line with prior works [20, 19, 42], we sample 2048 points from each 3D shape in segmentation tasks, resulting 128 point patches in the masked autoencoding pipeline (see Sec. 3).
Table 3 reports the mean Intersection over Union (mIoU) across all instances, and specific IoU values for each category on the ShapeNetPart benchmark. Point-MAE-Zero performs comparably to Point-MAE-SN, despite the lack of semantic information in the procedurally generated 3D shapes. This result suggests the absence of semantic input does not significantly hinder the learning of 3D point cloud representations using the masked autoencoding self-supervised learning framework.
Consistent with our findings in the object classification task in Sec. 4.1, both Point-MAE-Zero and Point-MAE-SN outperform models trained from scratch and prior works that use alternative network architectures[20, 21, 28] or pretraining methods [27, 42].
@@
4.3 Masked Point Cloud Completion
The goal of masked point cloud completion is to reconstruct masked points in input 3D point clouds, serving as a pretext task for learning 3D representations within the masked autoencoding framework [19] (see Fig. 1 and Sec. 3).
During pretraining, points in the input are grouped into patches, with a certain proportion of these patches randomly masked (e.g., 60% by default). Only the visible patches are passed to the encoder, while the centers of the masked patches are provided to the decoder as guidance for the masked autoencoding task. After pretraining, the network can reconstruct missing points in an input point cloud, even without the guidance of masked patch centers.
In our experiments, we test both settings — with guidance and without guidance — using pretrained Point-MAE-SN and Point-MAE-Zero. We evaluate performance on the ShapeNet test split and 2000 procedurally generated 3D shapes, neither of which was used during pretraining.
Fig. 2 presents qualitative comparisons between Point-MAE-SN and Point-MAE-Zero. Surprisingly, despite being trained exclusively on procedurally generated 3D shapes without semantic content, Point-MAE-Zero can still predict masked points on recognizable 3D objects (e.g., airplanes and chairs from ShapeNet) even without guidance points. We hypothesize that both models leverage symmetry to estimate missing parts (e.g., wings on an airplane or legs on a chair). Notably, Point-MAE-SN also performs well in predicting missing points on synthetic point clouds that were unseen in its pretraining dataset (i.e., ShapeNet).
Tab. 4 provides a quantitative comparison. Point-MAE-Zero and Point-MAE-SN exhibit very similar performance across both ShapeNet and synthetic shapes, though each model performs slightly better on its respective in-domain data. Performance for both models declines significantly when guidance points are removed from the decoder.
In summary, Fig. 2 and Tab. 4 highlight that the representations learned through the masked autoencoding strategy primarily capture geometric features rather than high-level semantic features during pretraining.
@@ 4.3 Masked Point Cloud Completion
Table 4: Masked Point Cloud Completion. The table reports the $L_{2}$ Chamfer distance (lower is better) between predicted masked points and ground truth on the test set of ShapeNet and procedurally synthesized 3D shapes. With Guidance: center points of masked patches are added to mask tokens in the pretrained decoder, guiding masked point prediction during Point-MAE training. Without Guidance: no information from masked patches is available during training.
4.4 Analysis
@@
Complexity of Synthetic 3D shapes.
We study how the geometric complexity of synthetic dataset contributes to performance in pretraining and downstream tasks. Specifically, we consider the following configurations, each with progressively increasing complexity:
  • (a) Single Primitive: Objects consist of a single primitive shape, each undergoing individual affine transformations.
  • (b) Multiple Primitives (up to 3): Objects contain up to three primitive shapes combined.
  • (c) Complex Primitives (up to 9): Objects are formed from up to nine primitive shapes.
  • (d) Shape Augmentation: Shapes are further modified through boolean difference operations and wire-frame conversions.
@@
Dataset Size.
Fig. 4 illustrates the effect of dataset size (i.e., the number of procedurally generated 3D shapes) on Point-MAE-Zero’s performance in the shape classification task on the PB-T50-RS benchmark.
@@
Figure 6: t-SNE visualization of 3D shape representations. (a) displays the distribution of representations from transformer encoders with different initialization strategies: randomly initialized (Scratch), pre-trained with Point-MAE on ShapeNet (Point-MAE-SN), and pre-trained on procedurally generated 3D shapes (Point-MAE-Zero). (b) shows the distribution of shape representations after fine-tuning on the target tasks, which include object classification on ModelNet40 (top row) and ScanObjectNN (bottom row). Each point represents a 3D shape while the color denotes the semantic categories.
Efficiency of Transfer Learning.
Fig. 5 shows the learning curves for training from scratch, Point-MAE-SN, and Point-MAE-Zero on shape classification tasks in the transfer learning setting. Both Point-MAE-SN and Point-MAE-Zero demonstrate faster training convergence and higher test accuracy compared to training from scratch. This trend is consistent across both ModelNet40 and ScanObjectNN benchmarks.
t-SNE Visualization.
Fig. 6 visualizes the distribution of 3D shape representations from Point-MAE-SN and Point-MAE-Zero via t-SNE [26], before and after fine-tuning on specific downstream tasks. It also includes representations from a randomly initialized neural network as a reference.
@@ 5 Discussion
In this work, we propose to learn 3D representations from synthetic data automatically generated using procedural 3D programs. We conduct an empirical analysis of Point-MAE [19], a state-of-the-art self-supervised learning method, and perform extensive comparisons with learning from well-curated, semantically meaningful 3D datasets.
We demonstrate that learning with procedural 3D programs performs comparably to learning from recognizable 3D models, despite the lack of semantic content in synthetic data. Our experiments highlights the importance of geometric complexity and dataset size in synthetic datasets for effective 3D representation learning. Our analysis further reveals that Point-MAE primarily learns geometric structures (e.g., symmetry) rather than high-level semantics.
@@ 6 Acknowledgement
We provide more implementation details, additional experimental results, and visualizations in this supplementary material. In Sec. A, we present our training-from-scratch baseline, which surpasses the results reported in Point-MAE [19]. In Sec. B, we introduce a more rigorous evaluation protocol using a dedicated validation set instead of relying on the test set for validation. Moreover, in Sec. C, we provide linear probing results on ModelNet40 and three ScanObjectNN variants to compare the performance of Point-MAE-SN and Point-MAE-Zero. Lastly, we provide additional visualization in Sec. D.

### v2

@@ 4.1 Object Classification
OcCo [37] 92.1 84.9 85.5 78.8
Point-BERT [55] 93.2 87.4 88.1 83.1
Point-MAE-Scratch [55] 91.4 79.9 80.6 77.2
Point-MAE-SN [26] 93.8 90.0 88.3 85.2
Point-MAE-Zero 93.0 90.4 88.6 85.5
Point-M2AE-Scratch 92.2 90.0 87.6 85.6
Point-M2AE-SN [59] 94.0 91.2 88.8 86.4
Point-M2AE-Zero 92.9 90.4 89.8 87.0
PCP-MAE-Scratch 91.5 88.8 88.5 83.8
PCP-MAE-SN [61] 94.0 95.5 94.3 90.4
PCP-MAE-Zero 92.4 94.0 92.3 90.5
@@
Table 1 summarizes object classification results across several settings. On ModelNet40, the “-Zero” variants (e.g., Point-MAE-Zero, Point-M2AE-Zero, PCP-MAE-Zero) generally fall slightly behind their ShapeNet‐pretrained counterparts (“-SN”), reflecting the larger domain gap between synthetic shapes and the clean 3D models in ModelNet40. By contrast, on ScanObjectNN—which contains real-world scans with broader geometric variability—the “-Zero” models often match or exceed the performance of their “-SN” counterparts. For instance, PCP-MAE-Zero outperforms PCP-MAE-SN on the PB-T50-RS variant, and Point-M2AE-Zero closely matches or exceeds Point-M2AE-SN in several cases. These findings indicate that the diverse geometry in procedurally synthesized data can be advantageous for certain real-world tasks. Meanwhile, all pretrained models (including both “-SN” and “-Zero”) surpass their respective from-scratch baselines and outperform existing self-supervised approaches (e.g., OcCo [37] and Point-BERT [55]).
@@
Point-M2AE-Scratch 87.5$\pm$2.6 90.0$\pm$5.5 86.4$\pm$3.2 89.6$\pm$4.3
Point-M2AE-SN [26] 93.4$\pm$3.1 96.2$\pm$1.5 91.8$\pm$4.5 92.9±$\pm 3.2$
Point-M2AE-Zero 91.4$\pm$1.8 94.2$\pm$2.2 88.3$\pm$3.7 91.0$\pm$2.9
PCP-MAE-Scratch 86.4$\pm$2.6 85.0$\pm$6.0 88.9$\pm$4.1 90.7$\pm$4.2
PCP-MAE-SN [61] 97.4$\pm$2.3 99.1$\pm$0.8 93.5$\pm$3.7 95.9$\pm$2.7
PCP-MAE-Zero 95.3$\pm$3.4 98.4$\pm$1.4 91.5$\pm$4.4 94.7$\pm$3.0
Table 2: Few-shot classification on ModelNet40. We evaluate performance on four $n$-way, $m$-shot configurations. For example, 5w/10s denotes a 5-way, 10-shot classification task. The table reports the mean classification accuracy (%) and standard deviation across 10 independent runs for each configuration. Top: Results from existing methods for comparison. Bottom: Comparison with our baseline methods. Note that results for Point-M2AE-SN are reproduced using publicly available code with our own configuration, as the original configuration was not provided.
@@
We evaluate few-shot classification on ModelNet40 using standard $n$-way, $m$-shot protocols, where $n$ denotes the number of randomly selected classes and $m$ the number of examples per class. Each evaluation samples 20 unseen instances from each class. We repeat this procedure 10 times, reporting mean accuracy (%) and standard deviation. Table 2 presents results for $n=\{5,10\}$ and $m=\{10,20\}$. Similar to transfer learning experiments, Point-MAE-Zero performs slightly below Point-MAE-SN, likely due to the larger domain gap between procedural shapes and ModelNet40 data. Nevertheless, both methods significantly outperform scratch-trained models and previous approaches such as DGCNN [38] and Transformer-OcCo [37].
@@ 4.2 Part Segmentation
The 3D part segmentation task aims to assign a part label to each point in a shape. We evaluate our methods and baselines on ShapeNetPart [54], which contains 16,881 models across 16 object categories. Consistent with previous works [27, 26, 55], we sample 2,048 points from each shape, resulting in 128 patches in our masked autoencoding pipeline (see Sec. 3).
Table 3 presents the mean Intersection-over-Union (mIoU) across all instances, along with per-category IoU. Across various models, both Point-MAE-Zero and Point-MAE-SN deliver comparable performance, indicating that procedurally generated shapes can learn robust 3D representations without explicit semantic content. Similarly, Point-M2AE-Zero and PCP-MAE-Zero achieve results on par with their ShapeNet-pretrained counterparts, further highlighting the versatility of procedural data in self-supervised representation learning.
In line with our observations in Sec. 4.1, the “-Zero” and “-SN” models surpass scratch-trained baselines and earlier methods that use different architectures [27, 28, 38] or alternative pretraining strategies [37, 55]. Despite lacking high-level semantic cues, these procedurally trained autoencoders still capture sufficient geometric structure to achieve strong segmentation performance.
@@
Point-M2AE-Scratch 84.7 85.1 86.8 88.6 81.1 91.5 79.9 92.1 87.8
Point-M2AE-SN [59] 85.0 84.5 87.2 89.3 81.1 91.8 80.1 92.0 89.2
Point-M2AE-Zero 84.9 85.3 87.3 88.7 81.1 91.7 79.4 91.9 88.2
PCP-MAE-Scratch 83.8 84.3 83.1 88.7 80.3 91.2 77.1 92.0 88.1
PCP-MAE-SN [61] 84.3 85.0 84.0 88.7 81.0 91.6 77.6 91.8 87.6
PCP-MAE-Zero 84.4 84.6 84.3 88.5 81.7 91.5 81.1 92.1 87.0
@@
Point-M2AE-Scratch 85.7 96.0 76.4 95.4 85.5 63.8 76.3 82.4
Point-M2AE-SN [59] 86.4 95.8 77.7 95.3 85.2 65.3 77.0 82.2
Point-M2AE-Zero 85.8 96.2 76.6 94.9 84.8 64.4 76.8 82.5
PCP-MAE-Scratch 84.9 95.0 76.0 95.0 85.0 63.2 75.4 81.0
PCP-MAE-SN [61] 85.8 96.4 76.1 95.2 84.8 64.0 77.4 81.4
PCP-MAE-Zero 86.0 96.1 76.6 94.6 85.1 63.6 76.8 80.4
@@
Figure 3: Masked Point Cloud Completion. This figure visualizes shape completion results with Point-MAE-SN and Point-MAE-Zero on the ShapeNet test split and procedurally synthesized 3D shapes. Left: Ground truth point clouds and masked inputs (60% mask ratio). Middle: Completions guided by masked input patch centers [26]. Right: Reconstructions without any guidance points. The $L_{2}$ Chamfer distance (lower is better) between the predicted 3D point clouds and the ground truth is displayed below each reconstruction.
@@
Point-M2AE-SN [59] 0.002 0.005 0.007 0.011
Point-M2AE-Zero 0.003 0.005 0.010 0.009
PCP-MAE-SN [61] - - 0.016 0.028
PCP-MAE-Zero - - 0.016 0.028
Table 4: Masked Point Cloud Completion. The table reports the $L_{2}$ Chamfer distance (lower is better) between predicted masked points and ground truth on the test set of ShapeNet and procedurally synthesized 3D shapes. With Guidance: center points of masked patches are added to mask tokens in the pretrained decoder, guiding masked point prediction during inference. Without Guidance: Without Guidance: no information from masked patches is available during inference.
4.3 Masked Point Cloud Completion
The masked point cloud completion task aims to reconstruct masked regions of input 3D point clouds, serving as a self-supervised pretext for learning 3D representations [26] (see Fig. 2 and Sec. 3).
During pretraining, points are grouped into patches, with a subset of patches (60% by default) randomly masked. Only visible patches are encoded, while masked patch centers can optionally guide the decoder (“with guidance”) or be omitted entirely (“without guidance”). After pretraining, models can reconstruct masked points even without such guidance. We quantitatively compare Point-MAE and Point-M2AE pretrained on ShapeNet (“-SN”) and procedural shapes (“-Zero”) in both guidance conditions, using the ShapeNet test split and 2,000 unseen synthetic shapes (see Tab. 4). All methods perform slightly better on their in-domain data. Removing guidance significantly decreases performance across all methods, highlighting its importance during masked reconstruction. Notably, Point-MAE-Zero and Point-M2AE-Zero closely match or even surpass their SN counterparts in reconstructing synthetic shapes, and remain competitive on ShapeNet shapes despite the lack of semantic training signals. PCP-MAE is a special case since it predicts centers before decoding point cloud and we find PCP-MAE-SN and PCP-MAE-Zero achieve similar performances both on seen and unseen domains.
Fig. 3 further illustrates that procedural-only models (e.g., Point-MAE-Zero) effectively reconstruct familiar ShapeNet objects (e.g., airplane wings, chair legs) without semantic supervision, likely by exploiting geometric symmetries. Similarly, SN-pretrained models generalize effectively to synthetic shapes not encountered during pretraining. Overall, these findings from Fig. 3 and Tab. 4 underscore that masked autoencoding primarily captures geometric rather than semantic information, enabling robust reconstruction across domains.
Semantic Segmentation Instance Segmentation
Methods mIoU mAcc allAcc mAP AP50 AP25
MSC-scan [41] 73.85 81.80 90.49 39.75 60.51 76.49
MSC-Zero (1k) 72.69 80.80 90.13 39.03 58.57 75.24
MSC-Zero (2k) 73.86 82.03 91.18 40.81 62.28 76.26
Table 5: Masked Scene Contrast Results. Performance comparison between MSC-Scan and MSC-Zero with different amounts of pretraining data for semantic and instance segmentation tasks.
4.4 Scene-level 3D Understanding Tasks
Given the effectiveness of procedural 3D programs for pretraining self-supervised learning (SSL) models on 3D objects, a natural question arises: Can procedural 3D programs similarly benefit SSL for 3D scenes? We adopt Masked Scene Contrast (MSC) [41], a state-of-the-art SSL method for 3D scenes. We pretrain MSC on ScanNet [7] (1K scenes), commonly used for 3D scene SSL pretraining, and compare it against MSC pretrained on our procedurally generated scenes (denoted MSC-Zero). We conduct experiments with MSC-Zero using varying amounts of data (1K and 2K procedural scenes). MSC-Zero trained with 2K procedurally generated 3D scenes achieves outperforms MSC pretrained on ScanNet in both 3D semantic and instance segmentation tasks. This aligns with our observations on object-level 3D understanding tasks. We discuss the exact procedures of generating such data and implementation details in the supplementary materials.
@@ 4.4 Scene-level 3D Understanding Tasks
4.5 Analysis
Complexity of Synthetic 3D shapes. We examine how the geometric complexity of synthetic datasets impacts pretraining and downstream performance. We consider four progressively complex configurations: (a) Single Primitive: a single shape with affine transformations; (b) Multiple Primitives ($\leq$3): up to three combined shapes; (c) Complex Primitives ($\leq$9): up to nine combined shapes; (d) Shape Augmentation: further modified via boolean differences and wire-frame conversions.
@@ 4.5 Analysis
Figure 7: t-SNE visualization of 3D shape representations. (a) Shows representations from transformer encoders: Scratch, Point-MAE-SN (ShapeNet), and Point-MAE-Zero (procedural shapes). (b) Displays fine-tuned representations for object classification on ModelNet40 (top) and ScanObjectNN (bottom). Each point represents a 3D shape while the color denotes the semantic categories.
Dataset Size. Fig. 5 illustrates the effect of dataset size (i.e., the number of procedurally generated 3D shapes) on Point-MAE-Zero’s performance in the shape classification task on the PB-T50-RS benchmark.
@@
Efficiency of Transfer Learning. Fig. 6 shows the learning curves for training from scratch, Point-MAE-SN, and Point-MAE-Zero on shape classification tasks in the transfer learning setting. Both Point-MAE-SN and Point-MAE-Zero demonstrate faster training convergence and higher test accuracy compared to training from scratch. This trend is consistent across both ModelNet40 and ScanObjectNN benchmarks.
t-SNE Visualization. Fig. 7 visualizes the distribution of 3D shape representations from Point-MAE-SN and Point-MAE-Zero via t-SNE [35], before and after fine-tuning on specific downstream tasks. It also includes representations from a randomly initialized neural network as a reference.
@@ 5 Discussion
In this work, we propose to learn 3D representations from synthetic data automatically generated using procedural 3D programs. We conduct an comprehensive empirical analysis of existing 3D SSLs and perform extensive comparisons with learning from well-curated, semantically meaningful 3D datasets.
We demonstrate that learning with procedural 3D programs performs comparably to learning from recognizable 3D models, despite the lack of semantic content in synthetic data. Our experiments highlights the importance of geometric complexity and dataset size in synthetic datasets for effective 3D representation learning. Our analysis further reveals that existing 3D SSLs primarily learns geometric structures (e.g., symmetry) rather than high-level semantics.
@@ 6 Acknowledgement
We provide more implementation details, additional experimental results, and visualizations in this supplementary material. In Sec. A, we present our training-from-scratch baseline, which surpasses the results reported in Point-MAE [26]. In Sec. B, we introduce a more rigorous evaluation protocol using a dedicated validation set instead of relying on the test set for validation. Moreover, in Sec. C, we provide linear probing results on ModelNet40 and three ScanObjectNN variants to compare the performance of Point-MAE-SN and Point-MAE-Zero. Additional visualizations are provided in Sec.D. Details on generating scene-level procedural 3D programs are in Sec.E. Lastly, implementation details for 3D object SSLs and 3D scene SSLs are presented in Sec. F.
@@ D.2 More t-SNE Visualizations
Appendix E Details on Generating Scene-Level Procedural 3D Programs
We follow MegaSynth’s  [17] generation pipeline. First, the pipeline generates a scene floor plan. Each 3D box represents a different shape, with distinct colors indicating various object types. This defines the spatial structure of the scene. Next, we creates 3D objects by combining primitive shapes such as cubes and spheres. These objects undergo geometry augmentations, including scaling, deformation, and boolean operations, to introduce variations. Since we do not need texture and lighting, we omit the rest of the procedurals. Instead we sample points from each shape and randomly sample points from the layout in order to represent walls, ceilings and floors. We will release code to further assist the reproducibility of our results.
Appendix F Implementation details.
We closely follow each baseline’s open‐source configuration. Similar to prior work for 3D shape pretraining, we sample each point cloud to $p=1024$ points and divide it into $n=64$ patches, with each patch containing $k=32$ points via the KNN algorithm. The autoencoder consists of an encoder with 12 Transformer blocks and a decoder with 4 Transformer blocks, each block having a 384‐dimensional hidden size and 6 attention heads. During pretraining, we randomly sample 1024 points per shape and apply standard random scaling and translation. We train for 300 epochs using the AdamW optimizer [24] with a cosine decay schedule [23], an initial learning rate of 0.001, weight decay of 0.05, and a batch size of 128.
For scene-level SSL MSC, we use SparseUNet34 as the backbone with a hierarchical encoder-decoder structure. The encoder consists of depths $[2,3,4,6]$ and channels $[32,64,128,256]$, while the decoder follows depths $[2,2,2]$ with channels $[256,128,64,64]$. A kernel size of 3 is used in both encoding and decoding, with a pooling stride of $[2,2,2,2]$. The pre-training phase employs SGD [31] with a cosine decay scheduler [23], an initial learning rate of 0.1, a weight decay of $1e^{-4}$, and a momentum of 0.8. We use a batch size of $32$ and train on ScanNet for 600 epochs, with 6 warmup epochs. For fine-tuning, we use a similar SGD [31] with cosine decay setup [23], but with a learning rate of 0.05, weight decay of $1e^{-4}$, and momentum of 0.9. The batch size is 48, with 40 warmup epochs, and the model is trained for 600 epochs. We will release code to further assist the reproducibility of our results.